The Blue Company makes boats. Job 589 is to build 10 boats. The company has 2 departments – assembly and fabrication. The assembly department allocates manufacturing overhead to jobs based on direct labor hours. The fabrication department allocates manufacturing overhead to jobs based on machine hours.

Please use the following information about the Blue Company;

                                                                                                    Assembly Department     Fabrication Department

Total fixed manufacturing overhead cost estimated

for all jobs during the whole year                                                           $480,000                         $560,000

Estimated direct labor hours for all jobs during the whole year          80,000                             60,000    

Estimated machine hours for all jobs during the whole year                20,000                             70,000

Estimated variable manufacturing overhead cost per direct labor hour   $9                                  ------

Estimated variable manufacturing overhead cost per machine hour      -----                                    $6                                                                                                                      

                                          

The following data is for Job 589:                      

                                                                                          Assembly Department     Fabrication Department

Direct materials cost                                                                       $400                                 $600

Direct labor cost                                                                              $900                                 $300

Direct labor hours used                                                                    70                                      20

Machine hours used                                                                         30                                       50

Please calculate for Job 589:

a. the total direct materials cost

b. the total direct labor cost

c. The predetermined overhead rate for the Assembly department

d. The predetermined overhead rate for the Fabrication department

e. The manufacturing overhead cost applied to job 589 by the Assembly department

f. The manufacturing overhead cost applied to job 589 by the Fabrication department

g. The total manufacturing overhead cost applied to job 589

h. The total cost of job 589

i. The cost per unit (per boat)
Answer #1

a) Total direct material cost = 400+600 = 1000

b) Total direct labor cost = 900+300 = 1200

c) Predetermine overhead rate for the assembly = (480000/80000+9) = 15 per DLH

d) Predetermine overhead rate for the Fabrication = (560000/70000+6) = 14 per MH

e) Total manufacturing overhead cost applied to Job 589 by the Assembly department = 70*15 = $1050

f) Total manufacturing overhead cost applied to Job 589 by the Fabrication department = 50*14 = 700

g) Total manufacturing overhead cost applied to Job 589 = 1050+700 = $1750

h) Total cost of Job 589 = 1000+1200+1750 = 3950

i) The cost per unit = 3950/10 = $395